### v1

@@ Abstract
chiappori2025changes study several indices that measure the assortativeness of a matching, including the aggregate likelihood ratio. We provide a counterexample showing that their axiomatization of the aggregate likelihood ratio is not valid as stated. We identify the exact class of indices characterized by the axioms in chiappori2025changes. We then show that the axiomatization of the aggregate likelihood ratio can be recovered by adding new axioms.
@@ 1 Introduction
The degree of assortative matching between men and women—how similar partners are in terms of education, age, income, and other socioeconomic characteristics—has long been viewed as a central determinant of inequality and welfare in modern economies (chiappori2023mating). If high-income or highly educated individuals increasingly marry one another, household income inequality may rise even when individual wage distributions remain stable. Understanding the extent and evolution of assortative matching is therefore crucial for interpreting long-run trends in inequality and social mobility.
In a pioneering study, chiappori2025changes propose an axiomatic foundation for indices of assortative matching and provide empirical evidence on changes in marital sorting in the United States. Their framework offers a valuable basis for interpreting trends in marital sorting in subsequent work. In this note, we show that the axiomatization of one measure, the aggregate likelihood ratio, is not correct as currently stated.111eika2019educational define a type-specific likelihood ratio as the ratio of the observed probability that a man and a woman of the same type (e.g., age) marry to the corresponding probability under random matching. The aggregate likelihood ratio is the weighted average of these type-specific likelihood ratios, with weights given by the expected mass of like-type couples under random matching. The measure has been applied to sociology and demographic studies such as pesando2021educational. We identify the exact class of indices characterized by the axioms in chiappori2025changes (Theorem 1). We then propose a refinement that corrects the statement and restores the intended result, while preserving the spirit and practical usefulness of their original approach (Theorem 2).
@@ 2 Model
Our notation follows that of chiappori2025changes. We consider two-type markets, i.e., there are two types of men and two types of women. Let $\mathcal{M}=\mathbb{R}_{\geq 0}^{4}\backslash\{\mathbf{0}\}$ denote the entire collection of possible matching patterns. We represent a matching pattern $M\in\mathcal{M}$ by a $2\times 2$ matrix given by
@@
An index is a function $I:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$. For each $M\in\mathcal{M}$, $I(M)$ represents how assortative the matching matrix is. Among several indices considered in chiappori2025changes, we focus on the following one.
@@
We revisit the axioms in chiappori2025changes. Consider a domain $\mathcal{M}^{\prime}\subseteq\mathcal{M}$ with $\mathcal{M}^{\prime}\neq\emptyset$. We define axioms for an index $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$. [^2] We note that chiappori2025changes’s (chiappori2025changes) define axioms for an index $I:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$. The following three axioms are called “invariance axioms”.
@@
To axiomatize ALR, chiappori2025changes introduced the following axiom.
@@
We reproduce Theorem 3 of chiappori2025changes.
@@ 4 Recovering axiomatization of the aggregate likelihood ratio
In this section, we recover the axiomatization of ALR in chiappori2025changes’s (chiappori2025changes). To this end, we first identify the exact class of indices that satisfy the axioms in Theorem CDMZ3.
@@
Theorem 1 suggests that, to pin down the class of positive multiples of ALR, we need additional axioms that force $\beta=0$ in (7). We consider an axiom that is interpreted as the opposite of the maximum homogamy axiom in chiappori2025changes (defined in Section 5). As in Section 2, we define axioms for an index $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$, where $\mathcal{M}^{\prime}\subseteq\mathcal{M}$, $\mathcal{M}^{\prime}\neq\emptyset$.
  • Maximum Heterogamy. For each $M\in\mathcal{M}^{\prime}$ and $(a,b,c,d)\in\mathcal{M}^{\prime}$ such that $b>0,c>0,a=d=0$, we have $I(M)\geq I(a,b,c,d)$.
@@
Proof of the “only if” direction: By Theorem 1, there exist $\alpha,\beta\in\mathbb{R}$ with $\alpha>\beta\geq 0$ such that $I$ is given as in (7) on the domain $\mathcal{M}_{>0}$. If $I$ satisfies maximum heterogamy, then $\beta=0$. Indeed, if $\beta>0$, by letting $b,c\rightarrow+\infty$ while keeping $a=d=0$, the value of $I(a,b,c,d)$ grows arbitrarily large, contradicting maximum heterogamy.
@@
Note that for any $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, there exists a sequence $\{M^{k}\}_{k=1}^{\infty}\subseteq\mathcal{M}_{>0}$ such that $M^{k}\rightarrow M$. Therefore, by continuity, the value of $I(M)$ for $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$ is also given by (17). ∎
5 Concluding remarks
@@
We have shown that Theorem 3 of chiappori2025changes does not hold as currently stated and we proposed a modification. We note that Theorem 2 of chiappori2025changes does not hold as currently stated either. Specifically, chiappori2025changes consider the following index called the normalized trace:
@@
The following is Theorem 2 of chiappori2025changes.
@@
  • Maximum Homogamy. For each $M\in\hat{\mathcal{M}}$ and $(a,b,c,d)\in\hat{\mathcal{M}}$ such that $ad>0$ and $bc=0$, we have $I(a,b,c,d)\geq I(M)$.
@@
One can verify that this modified index satisfies the three invariance axioms, marginal monotonicity, maximum homogamy, and population decomposability. [^4] This modified index also satisfies other two monotonicity axioms in chiappori2025changes, namely diagonal monotonicity and off-diagonal monotonicity. However, this modified index is not a positive affine transformation of the normalized trace. To see this, let $M=(a,b,c,d)\in\hat{\mathcal{M}}$. We consider two cases.

### v2

@@ Abstract
Chiappori et al. (2025) study several indices of assortativeness in matching, including the aggregate likelihood ratio and the odds ratio. We provide a counterexample showing that their axiomatization of the aggregate likelihood ratio is not valid as stated. We identify the exact class of indices characterized by the axioms in Chiappori et al. (2025). We then show that the axiomatization of the aggregate likelihood ratio can be recovered by adding new axioms. In addition, we point out errors in the axiomatizations of other measures in Chiappori et al. (2025). Finally, we offer a generalization of the odds ratio from two-type markets to multi-type markets.
@@ 1 Introduction
The degree of assortative matching between men and women—how similar partners are in terms of education, age, income, and other socioeconomic characteristics—has long been viewed as a central determinant of inequality and welfare in modern economies (Chiappori and Salanié 2023). If high-income or highly educated individuals increasingly marry one another, household income inequality may rise even when individual wage distributions remain stable. Understanding the extent and evolution of assortative matching is therefore crucial for interpreting long-run trends in inequality and social mobility.
In a pioneering study, Chiappori et al. (2025) propose an axiomatic foundation for indices of assortative matching and provide empirical evidence on changes in marital sorting in the United States. Their framework offers a valuable basis for interpreting trends in marital sorting in subsequent work. In this note, we show that the axiomatization of one measure, the aggregate likelihood ratio, is not correct as currently stated. [^1] Eika et al. (2019) define a type-specific likelihood ratio as the ratio of the observed probability that a man and a woman of the same type (e.g., age) marry to the corresponding probability under random matching. The aggregate likelihood ratio is the weighted average of these type-specific likelihood ratios, with weights given by the expected mass of like-type couples under random matching. The measure has been applied to sociology and demographic studies such as Pesando (2021). We identify the exact class of indices characterized by the axioms in Chiappori et al. (2025) (Theorem 1). We then propose a refinement that corrects the statement and restores the intended result, while preserving the spirit and practical usefulness of their original approach (Theorem 2).
Chiappori et al. (2025) axiomatize two measures other than the aggregate likelihood ratio: the odds ratio and the normalized trace. We show that the axiomatizations of these masures are not correct either. In the final section, we discuss a generalization of the odds ratio from two-type markets to multi-type markets.
@@ 2 Model
Our notation follows that of Chiappori et al. (2025). We consider two-type markets, i.e., there are two types of men and two types of women. Let $\mathcal{M}=\mathbb{R}_{\geq 0}^{4}\backslash\{\mathbf{0}\}$ denote the entire collection of possible matching patterns. We represent a matching pattern $M\in\mathcal{M}$ by a $2\times 2$ matrix given by
@@
An index is a function $I:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$. For each $M\in\mathcal{M}$, $I(M)$ represents how assortative the matching matrix is. Among several indices considered in Chiappori et al. (2025), we focus on the following one.
@@
We revisit the axioms in Chiappori et al. (2025). Consider a domain $\mathcal{M}^{\prime}\subseteq\mathcal{M}$ with $\mathcal{M}^{\prime}\neq\emptyset$. We define axioms for an index $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$. [^2] We note that Chiappori et al.’s (2025) define axioms for an index $I:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$. The following three axioms are called “invariance axioms”.
@@
To axiomatize ALR, Chiappori et al. (2025) introduced the following axiom.
@@
We reproduce Theorem 3 of Chiappori et al. (2025).
@@ 4 Recovering axiomatization of the aggregate likelihood ratio
In this section, we recover the axiomatization of ALR in Chiappori et al. (2025). To this end, we first identify the exact class of indices that satisfy the axioms in Theorem CDMZ3.
@@
Theorem 1 suggests that, to pin down the class of positive multiples of ALR, we need additional axioms that force $\beta=0$ in (7). We consider an axiom that is interpreted as the opposite of the maximum homogamy axiom in Chiappori et al. (2025) (defined in Section 5). As in Section 2, we define axioms for an index $I:\mathcal{M}^{\prime}\rightarrow\mathbb{R}_{\geq 0}$, where $\mathcal{M}^{\prime}\subseteq\mathcal{M}$, $\mathcal{M}^{\prime}\neq\emptyset$.
  • Maximum Heterogamy. For each $M\in\mathcal{M}_{>0}$ and $(a,b,c,d)\in\mathcal{M}^{\prime}$ such that $b>0,c>0,a=d=0$, we have $I(M)\geq I(a,b,c,d)$.
@@
Proof of the “only if” direction: By Theorem 1, there exist $\alpha,\beta\in\mathbb{R}$ with $\alpha>\beta\geq 0$ such that $I$ is given as in (7) on the domain $\mathcal{M}_{>0}$. Note that for any $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$, there exists a sequence $\{M^{k}\}_{k=1}^{\infty}\subseteq\mathcal{M}_{>0}$ such that $M^{k}\rightarrow M$. Therefore, by continuity, the value of $I(M)$ for $M\in\tilde{\mathcal{M}}\setminus\mathcal{M}_{>0}$ is also given by (17).
If $I$ satisfies maximum heterogamy, then $\beta=0$. Indeed, if $\beta>0$, by letting $b,c\rightarrow+\infty$ while keeping $a=d=0$, the value of $I(a,b,c,d)$ grows arbitrarily large, contradicting maximum heterogamy. ∎
5 Remarks on axiomatizations of other measures
@@
We have shown that Theorem 3 of Chiappori et al. (2025) does not hold as currently stated and we proposed a modification. We note that Theorems 1 and 2 of Chiappori et al. (2025) do not hold as currently stated either.
5.1 Odds ratio
For $(a,b,c,d)\in\mathcal{M}$, we define the odds ratio by
$I_{O}(M)=\left\{\begin{array}[]{l}\frac{ad}{bc}\text{ if }b\neq 0\text{ and }c\neq 0,\\ +\infty\text{ if }bc=0.\end{array}\right$.
A preorder $\succeq$ is a reflexive and transitive binary relation on $\mathcal{M}$. We revisit Chiappori et al.’s (2025) axioms on a preorder $\succeq$.
  • Marginal Independence. For any $M\in\mathcal{M}$ and $\lambda>0$,
    $a$ $b$ $c$ $d$ $\sim$ $\lambda a$ $\lambda b$ $c$ $d$ $\sim$ $a$ $b$ $\lambda c$ $\lambda d$ $\sim$ $\lambda a$ $b$ $\lambda c$ $d$ $\sim$ $a$ $\lambda b$ $c$ $\lambda d$ .
  • Maximum Homogamy. For any $M\in\mathcal{M},ad>0$, and $bc=0,(a,b,c,d)\succeq M$.
  • Marginal Monotonicity (for preorders). For every pair of positive matching matrices $M=(a,b,c,d)\in\mathcal{M}^{\prime}$ and $M^{\prime}=\left(a^{\prime},b^{\prime},c^{\prime},d^{\prime}\right)$ with the same marginal distributions (i.e., $a+c=a^{\prime}+c^{\prime},a+b=a^{\prime}+b^{\prime},d+b=d^{\prime}+b^{\prime}$, and $\left.d+c=d^{\prime}+c^{\prime}\right)$, $M\succ M^{\prime}$ if and only if $a>a^{\prime}$ (equivalently, $b<b^{\prime},c<c^{\prime}$, or $\left.d>d^{\prime}\right)$.
The following is Theorem 1 of Chiappori et al. (2025).
Theorem CDMZ1.
The odds ratio induces the unique preorder that satisfies marginal monotonicity, marginal independence, and maximum homogamy.
To introduce a counterexample to this theorem, we define three subsets of $\mathcal{M}$.
  • $\mathcal{M}_{\infty}=\{(a,b,c,d)\mid bc=0,ad>0\}$.
  • $\widetilde{\mathcal{M}}_{\infty}=\{(a,b,c,d)\mid bc=0,ad=0\}$
  • $\mathcal{M}_{0}=\{(a,b,c,d)\mid bc>0\text{ and }ad=0\}$.
Then, it holds that $\mathcal{M}=\mathcal{M}_{\infty}\cup\widetilde{\mathcal{M}}_{\infty}\cup\mathcal{M}_{0}\cup\mathcal{M}_{>0}$. Note that $\mathcal{M}_{\infty}$ and $\widetilde{\mathcal{M}}_{\infty}$ denote the set of matrices to which the odds ratio assigns $\infty$, $\mathcal{M}_{0}$ denotes the set of matrices to which the odds ratio assigns 0, and $\mathcal{M}_{>0}$ denotes the set of matrices to which the odds ratio assigns positive real number.
We define a preorder $\succeq^{*}$ as follows:
  • (1)
    for each $M\in\mathcal{M}_{\infty}$ and $M^{\prime}\in\widetilde{\mathcal{M}}_{\infty}\cup\mathcal{M}_{0}\cup\mathcal{M}_{>0}$, we have $M\succ^{*}M^{\prime}$.
    for each $M,M^{\prime}\in\mathcal{M}_{\infty}$, we have $M\sim^{*}M^{\prime}$.
  • (2)
    for each $M\in\widetilde{\mathcal{M}}_{\infty}$ and $M^{\prime}\in\mathcal{M}_{0}\cup\mathcal{M}_{>0}$, we have $M\succ^{*}M^{\prime}$.
    for each $M,M^{\prime}\in\widetilde{\mathcal{M}}_{\infty}$, we have $M\sim^{*}M^{\prime}$.
  • (3)
    for each $M\in\mathcal{M}_{0}$ and $M^{\prime}\in\mathcal{M}_{>0}$, we have $M\succ^{*}M^{\prime}$.
    for each $M,M^{\prime}\in\mathcal{M}_{0}$, we have $M\sim^{*}M^{\prime}$.
  • (4)
    for each $M,M^{\prime}\in\mathcal{M}_{>0}$, we have $M\succeq^{*}M^{\prime}$ if and only if $I_{O}(M)\geq I_{O}(M^{\prime})$.
Then, $\succeq^{*}$ satisfies all the three axioms of Theorem CDMZ1. First, $\succeq^{*}$ satisfies marginal monotonicity because this axiom bears only on matrices in $\mathcal{M}_{>0}$, and in this space, $\succeq^{*}$ is induced from the odds ratio. Moreover, $\succeq^{*}$ satisfies maximum homogamy by Condition (1). We show that $\succeq^{*}$ satisfies marginal independence.
  • For $M\in\mathcal{M}_{\infty}$ and $\lambda>0$, we have $\lambda M\in\mathcal{M}_{\infty}$. Therefore, by Condition (1), we have $M\sim^{*}\lambda M$.
  • For $M\in\widetilde{\mathcal{M}}_{\infty}$ and $\lambda>0$, we have $\lambda M\in\widetilde{\mathcal{M}}_{\infty}$. Therefore, by Condition (2), we have $M\sim^{*}\lambda M$.
  • For $M\in\mathcal{M}_{0}$ and $\lambda>0$, we have $\lambda M\in\mathcal{M}_{0}$. Therefore, by Condition (3), we have $M\sim^{*}\lambda M$.
  • For $M\in\mathcal{M}_{>0}$ and $\lambda>0$, we have $\lambda M\in\mathcal{M}_{>0}$. Since $\succ^{*}$ is induced from the odds ratio, we have $M\sim\lambda M$. Therefore, $\succeq^{*}$ satisfies marginal monotonicity.
We conclude that $\succeq^{*}$ satisfies all the three axioms. However, $\succeq^{*}$ is not induced from the odds ratio. To see this, consider the following two matrices.
[…]
Since $M\in\mathcal{M}_{\infty}$ and $M^{\prime}\in\widetilde{\mathcal{M}}_{\infty}$, Condition (1) implies $M\succ^{*}M^{\prime}$. However, according to the odds ratio,
$I_{O}(M)=+\infty=I_{O}(M^{\prime}).$
Therefore, $\succeq^{*}$ is not induced from the odds ratio.
One can also verify this negative result by considering the following two matrices.
[…]
Since $M\in\mathcal{M}_{0}$ and $M^{\prime}\in\mathcal{M}_{>0}$, Condition (3) implies $M\succ^{*}M^{\prime}$. However, according to the odds ratio,
$I_{O}(M)=0<1=I_{O}(M^{\prime}).$
Therefore, $\succeq^{*}$ is not induced from the odds ratio. We conclude that Theorem CDMZ1 is not correct.
Chiappori et al. (2025) are correct in showing that any preorder satisfying the axioms in Theorem CDMZ1 must coincide with the preorder induced by the odds ratio on the class of positive matching matrices $\mathcal{M}_{>0}$. However, this conclusion is not guaranteed on the remaining domain. To remedy this problem, we make two modifications. First, we strengthen Maximum Homogamy to eliminate the preference gap between matrices in $\mathcal{M}_{\infty}$ and $\widetilde{\mathcal{M}}_{\infty}$.
  • Maximum Homogamy+. For any $M\in\mathcal{M}$ and $(a,b,c,d)\in\mathcal{M}$ such that $bc=0$, we have $(a,b,c,d)\succeq M$.
The second modification is to impose Maximum Heterogamy as in Section 4 to gurantee that every matrix in $\mathcal{M}_{0}$ is evaluated as least assortative.
  • Maximum Heterogamy (for preorders). For each $M\in\mathcal{M}_{>0}$ and $(a,b,c,d)\in\mathcal{M}$ such that $b>0,c>0,a=d=0$ (i.e., $(a,b,c,d)\in\mathcal{M}_{0}$), we have $M\succeq(a,b,c,d)$.
The following theorem recovers the axiomatization of the odds ratio.
Theorem 4.
The odds ratio induces the unique preorder that satisfies marginal monotonicity, marginal independence, maximum homogamy+, and maximum heterogamy.
Note that maximum homogamy+ and maximum heterogamy uniquely pin down the ordinal ranking between a positive matching matrix and non-positive matching matrix, and between nonpositive matching matrices. [^4] Precisely, to prove $M\succ M^{\prime}$ for $M\in\mathcal{M}_{>0}$ and $M^{\prime}\in\mathcal{M}_{0}$, we need marginal monotonicity aside from maximum heterogamy. The proof proceeds as follows: by maximum heterogamy, we have $M\succeq M^{\prime}$. If $M\sim M^{\prime}$, then letting $M=(a,b,c,d)$ and $M_{\varepsilon}=(a-\varepsilon,b+\varepsilon,c+\varepsilon,d-\varepsilon)\in\mathcal{M}_{>0}$ for a sufficiently small $\varepsilon$, marginal monotonicity implies $M\succ M_{\varepsilon}$. But then, we have $M^{\prime}\sim M\succ M_{\varepsilon}$, contradicting maximum heterogamy. Since marginal monotonicity and marginal independence imply that the oreorder is induced from the odds ratio on the class of positive matching matrices, Theorem 4 holds.
5.2 Normalized trace
Specifically, Chiappori et al. (2025) consider the following index called the normalized trace:
@@ 5.2 Normalized trace
The following is Theorem 2 of Chiappori et al. (2025).
@@
  • Maximum Homogamy (for indices). For each $M\in\hat{\mathcal{M}}$ and $(a,b,c,d)\in\hat{\mathcal{M}}$ such that $ad>0$ and $bc=0$, we have $I(a,b,c,d)\geq I(M)$.
@@
One can verify that this modified index satisfies the three invariance axioms, marginal monotonicity, maximum homogamy, and population decomposability. [^5] This modified index also satisfies other two monotonicity axioms in Chiappori et al. (2025), namely diagonal monotonicity and off-diagonal monotonicity. However, this modified index is not a positive affine transformation of the normalized trace. To see this, let $M=(a,b,c,d)\in\hat{\mathcal{M}}$. We consider two cases.